# Van der Waals devices for surface-sensitive experiments

### v1

@@ IV Experimental Methods
Single crystals of 1T–TaS2 are grown by chemical vapor transport using I2 as a transport agent, following the method described in [4].
@@
Shadow mask fabrication: Stencils are fabricated from 250-µm thick Si wafers utilizing laser lithography and DRIE. The electrode design is exposed on a 9-µm thick film of positive tone photoresist (SPR 220 7.0, micro resist technology GmbH) using a direct-write laser lithography tool (DWL 66+, Heidelberg Instruments Mikrotechnik GmbH) at nm. The Si wafer is etched through in a DRIE tool (Omega Rapier 200-mm process module, SPTS Technologies Ltd.) using a variation of the highly anisotropic “Bosch process” optimized for low sidewall roughness and micron-scale features. It consists of alternating etching and passivation cycles. The combination of physical and chemical mechanisms enables highly directional etching through the entire wafer, while minimizing lateral etching and ensuring vertical sidewalls.
Our process is carried out at C with the electrostatic chuck biased at V and the coil current set to A. The sequence of three steps (see the process parameters listed in Tab. 2) starts with Si etching from the reaction between F radicals, from SF6 plasma, and Si that results in volatile SiF4 byproducts. Additional platen power is applied to accelerate plasma-generated ions toward the surface and remove material selectively. However, the chemical nature of this process leads to unwanted lateral etching. To counteract this effect, our process incorporates a passivation step, in which isotropic C4F8 plasma deposits a fluorocarbon polymer layer on both the surface and the sidewalls of the etched trenches. This is followed by an intermediate base etching step using a gaseous mixture of SF6 and O2. Platen bias power is applied to enhance anisotropy and remove most of the passivation layer at the base of the trenches, while preserving it on the sidewalls to protect them during the subsequent Si etch, which continues until the passivation layer is completely consumed. These steps are repeated cyclically, where each cycle (see Fig. 4) removes a controlled amount of Si and prevents lateral etching, enabling the formation of high-aspect-ratio structures.
Device design: The geometry of the shadow mask used for this proof-of-principle study, featuring a mmµm bridge, is designed to form two-terminal devices where exfoliated flakes cover the predefined contact gaps. The contact dimensions are optimized for the typical size and density of exfoliated flakes while ensuring compatibility with the spatial resolution of the micro-beam ARPES setup. A gap size of µm is found to provide a balance between the ability to resolve the area between contact pads and maintaining a high probability of flakes bridging. The vertical gap length (mm) determines the average amount of flakes covering the gap—longer gaps increase the probability of multiple flakes, while shorter ones are preferred for single-flake transport measurements (at the risk of having no flake bridging the gap). The large lateral extent of the contact pads to the sides (mm) ensures sufficient exposed area next to the tape loop for the fixation of wires to establish electrical contact to the device.
@@
Ti/Au contacts are deposited on the substrate using the shadow mask in an electron beam evaporator (BAK Uni, Evatec AG) at pressures below mbar. A 2-nm Ti layer is deposited first to promote adhesion between SiO2 and the 13-nm Au film [66]. The thickness of the Au layer is optimized to balance surface smoothness for exfoliation and mechanical durability during electrical pulsing.
@@
The devices are fabricated on 4-inch Si/SiO2 wafers (MicroChemicals GmbH) with a thickness of µm, p-type doping, a resistivity of mcm, bow/warp µm, and -nm of SiO2.
@@
After contact deposition, bulk 1T–TaS2 crystals with typical dimensions of mm in diameter and µm thickness are cleaved on tape loops, brought into contact with the substrate—aligning the bulk material by eye with the pre-patterned contacts—and gently pressed on to avoid trapping air pockets under the tape. Among other commonly used tapes, e.g. Nitto Blue Tape or Scotch Tape, we achieved the most reproducible results with silicone adhesive, UHV compatible Kapton tape.
The exfoliation process strongly depends on the cleanliness and smoothness of the Au surface as contamination and roughness diminish the adhesion between the vdW material and Au, leading to a lower exfoliation yield [21]. This occurs because airborne organic contaminants accumulate on the Au surface, turning it more hydrophobic and weakening the adhesion [27]. To minimize these effects, the transfer is performed rapidly after venting the evaporation chamber with N2. Material transfer within min consistently yields large-area flakes down to the monolayer limit, whereas longer time spans result in predominantly thicker (nm) flakes with lower yield.
@@
The Si wafer containing 24 devices is cleaved manually into dies and glued on standard flag-style copper plates (see Fig. S1) using Ag epoxy (EPO-TEK H20E, Epoxy Technology Inc.). Electrical connections to the device are also made using Ag epoxy by attaching insulated copper wires to the Au pads. The curing of the epoxy is performed under conditions that minimize thermal stress: We heat the fully assembled sample holder to the lowest temperature specified by the manufacturer that ensures full curing of the epoxy, i.e. C, thereby avoiding thermal expansion that could strain the vdW flake or introduce residues from melting the adhesive material of the tape.
@@
Electrical measurements are carried out using a pulsed current source (Keithley 6221, Tektronix, Inc.) and nanovoltmeter (Keithley 2182). Resistance values are obtained from a linear fit to the slope of IV curves on a µA current interval. Pulsed IV curves to measure the switching to the hidden state are recorded using a pulsed sweep mode, applying current pulses from 0 to 16 mA in µA steps. Each pulse has a width of µs, but the voltage is measured only during the second half of the pulse, i.e. after a µs source delay. Between pulses, a pause of ms minimizes heating effects.
@@
ARPES measurements are conducted at the Spectromicroscopy endstation at Elettra Sincrotrone Trieste with a photon energy of eV and an energy resolution of meV. A beam footprint of µm2 is achieved through focusing with a zone plate and a Schwarzschild objective. Spectra are acquired using a hemispherical analyzer (MB Scientific AB A-1) equipped with a 2D deflector system, accepting a cone. Devices are cleaved at room temperature in a preparation chamber (mbar pressures) and transferred to the manipulator head in the main chamber (mbar). The cooling rate is controlled at –K/min to cross the NC-CCDW transition of 1T–TaS2 in a controlled manner. The binding energy axis is aligned to by fitting the angle-integrated spectrum with a Fermi-Dirac distribution.
@@
To analyze the Ta 47/2 core-level spectra, we employ a multi-step fitting procedure. First, we perform spatial clustering using a -means algorithm [62] to group different regions of interest based on spectral similarity. Within each cluster, all angle-integrated spectra are averaged to obtain the cluster mean and standard deviation. A Shirley-type background [63] is subtracted to account for inelastic scattering contributions. The core-level spectra are fitted using three Voigt profiles, corresponding to the three in-equivalent Ta sites of the star-of-David pattern forming the CDW state. The Voigt function, defined as the convolution of a Lorentzian and a Gaussian profile, accounts for both the intrinsic lifetime broadening of the core-hole state (Lorentzian) and the instrumental resolution of the measurement setup (Gaussian). The initial peak positions were set to eV, eV, and eV, based on prior studies on bulk crystals [64], and allowed to vary within eV. The initial amplitude ratios were set to 1:6:6 (A:B:C), and the Gaussian and Lorentzian widths constrained to and , respectively. The best fit parameters were determined using least-squares optimization. Intensity ratios were established by integrating the peak areas.
@@
AFM measurements were conducted after the ARPES experiment using a Dimension 3100 system (Bruker Corp.) to establish the thickness and homogeneity of the 1T–TaS2 flakes.
Scanning electron microscopy: High-resolution images of the shadow mask were obtained using a scanning electron microscope (Regulus 8230, Hitachi Corp.).
V Author Contributions
@@
VI Data Availability
The data that support the findings of this study are available from the corresponding authors upon request.
VII Conflict of Interest
@@ VIII Acknowledgments
Programme, within Grant Agreement 810451 (HERO). R.V., A.M., Y.V and D.M. thank ARIS for funding the research program P1-0400, and D.M. thanks ARIS for funding the

### v2

@@ 4 Experimental methods
Single crystals of 1T–TaS2 are grown by chemical vapor transport using I2 as a transport agent, following previously established procedures 4.
Shadow mask fabrication:
Stencils are fabricated from 250-µm thick Si wafers utilizing laser lithography and DRIE. The electrode design is exposed on a 9-µm thick film of positive tone photoresist (SPR 220 7.0, micro resist technology) using a direct-write laser lithography tool (DWL 66+, Heidelberg Instruments Mikrotechnik) at nm. The Si wafer is etched through in a DRIE tool (Omega Rapier 200-mm process module, SPTS Technologies) using a variation of the highly anisotropic “Bosch process” optimized for low sidewall roughness and micron-scale features. It consists of alternating etching and passivation cycles. The combination of physical and chemical mechanisms enables highly directional etching through the entire wafer, while minimizing lateral etching and ensuring vertical sidewalls.
Our process is carried out at C with the electrostatic chuck biased at V and the coil current set to A. The sequence of three steps (see the process parameters listed in Tab. 2) starts with Si etching from the reaction between F radicals, from SF6 plasma, and Si that results in volatile SiF4 byproducts. Additional platen power is applied to accelerate plasma-generated ions toward the surface and remove material selectively. The chemical nature of this process leads to unwanted lateral etching. To counteract this effect, our process incorporates a passivation step in which isotropic C4F8 plasma deposits a fluorocarbon polymer layer on both the surface and the sidewalls of the etched trenches. This is followed by an intermediate base etching step using a gaseous mixture of SF6 and O2. Platen bias power is applied to enhance anisotropy and remove most of the passivation layer at the base of the trenches, while preserving it on the sidewalls to protect them during the subsequent Si etch which continues until the passivation layer is completely consumed. These steps are repeated cyclically, where each cycle (see Fig. 4) removes a controlled amount of Si and prevents lateral etching, enabling the formation of high-aspect-ratio structures.
@@
Device design:
The geometry of the shadow mask used for this proof-of-principle study, featuring a mmµm bridge, is designed to form two-terminal devices where exfoliated flakes cover the predefined contact gaps. The contact dimensions are optimized for the typical size and density of exfoliated flakes while ensuring compatibility with the spatial resolution of the micro-beam ARPES setup. A gap size of µm is found to provide a balance between the ability to resolve the area between contact pads and maintaining a high probability of flakes bridging. The vertical gap length (mm) determines the average amount of flakes covering the gap—longer gaps increase the probability of multiple flakes, while shorter ones are preferred for single-flake transport measurements (at the risk of having no flake bridging the gap). The large lateral extent of the contact pads to the sides (mm) ensures sufficient exposed area next to the tape loop for the fixation of wires to establish electrical contact to the device.
@@
Ti/Au contacts are deposited on the substrate using the shadow mask in an electron beam evaporator (BAK Uni, Evatec) at pressures below mbar. A 2-nm Ti layer is deposited first to promote adhesion between SiO2 and the 13-nm Au film 66. The thickness of the Au layer is optimized to balance surface smoothness for exfoliation and mechanical durability during electrical pulsing.
@@
The devices are fabricated on 4-inch Si/SiO2 wafers (MicroChemicals) with a thickness of µm, p-type doping, a resistivity of mcm, bow/warp µm, and nm of SiO2.
@@
After contact deposition, bulk 1T–TaS2 crystals with typical dimensions of mm in diameter and µm thickness are cleaved on tape loops, brought into contact with the substrate—aligning the bulk material by eye with the pre-patterned contacts—and gently pressed on to avoid trapping air pockets under the tape. Among other commonly used tapes, e.g. Nitto Blue Tape or Scotch Tape, we achieved the most reproducible results with silicone adhesive, UHV compatible Kapton tape. The exfoliation process strongly depends on the cleanliness and smoothness of the Au surface as contamination and roughness diminish the adhesion between the vdW material and Au, leading to a lower exfoliation yield 24. This occurs because airborne organic contaminants accumulate on the Au surface, turning it more hydrophobic and weakening the adhesion 27. To minimize these effects, the transfer is performed rapidly after venting the evaporation chamber with N2. Material transfer within min consistently yields large-area flakes down to the monolayer limit, whereas longer time spans result in predominantly thicker (nm) flakes with lower yield.
@@
The Si wafer containing 24 devices is cleaved manually into dies and glued on standard flag-style copper plates (see Fig. S1) using Ag epoxy (EPO-TEK H20E, Epoxy Technology). Electrical connections to the device are also made using Ag epoxy by attaching insulated copper wires to the Au pads. The curing of the epoxy is performed under conditions that minimize thermal stress: We heat the fully assembled sample holder to the lowest temperature specified that ensures full curing of the epoxy, i.e. C, thereby avoiding thermal expansion that could strain the vdW flake or introduce residues from melting the adhesive material of the tape.
@@
Measurements are carried out using a pulsed current source (Keithley 6221, Tektronix) and nanovoltmeter (Keithley 2182). Resistance values are obtained from a linear fit to the slope of IV curves on a µA current interval. Pulsed IV curves to measure the switching to the hidden state are recorded using a pulsed sweep mode, applying current pulses from 0 to 16 mA in µA steps. Each pulse has a width of µs, but the voltage is measured only during the second half of the pulse, i.e. after a µs source delay. Between pulses, a pause of ms minimizes heating.
@@
Measurements are conducted at the Spectromicroscopy endstation at Elettra Sincrotrone Trieste with a photon energy of eV and an energy resolution of meV. A beam footprint of µm2 is achieved through focusing with a zone plate and a Schwarzschild objective. Spectra are acquired using a hemispherical analyzer (MB Scientific AB A-1) equipped with a 2D deflector system, accepting a cone. Devices are cleaved at room temperature in a preparation chamber (mbar pressures) and transferred to the manipulator head in the main chamber (mbar). The cooling rate is controlled at –K/min to cross the NC-CCDW transition of 1T–TaS2 in a controlled manner. The binding energy axis is aligned to by fitting the angle-integrated spectrum with a Fermi-Dirac distribution.
@@
To analyze the Ta 47/2 core-level spectra, we employ a multi-step fitting procedure. First, we perform spatial clustering using a -means algorithm 62 to group different regions of interest based on spectral similarity. Within each cluster, all angle-integrated spectra are averaged to obtain the cluster mean and standard deviation. A Shirley-type background 63 is subtracted to account for inelastic scattering contributions. The core-level spectra are fitted using three Voigt profiles, corresponding to the three in-equivalent Ta sites of the star-of-David pattern forming the CDW state. The Voigt function, defined as the convolution of a Lorentzian and a Gaussian profile, accounts for both the intrinsic lifetime broadening of the core-hole state (Lorentzian) and the instrumental resolution of the measurement setup (Gaussian). The initial peak positions were set to eV, eV, and eV, based on prior studies on bulk crystals 64, and allowed to vary within eV. The initial amplitude ratios were set to 1:6:6 (A:B:C), and the Gaussian and Lorentzian widths constrained to and , respectively. The best fit parameters were determined using least-squares optimization. Intensity ratios were established by integrating the peak areas.
@@
Measurements are conducted after the ARPES experiment using a Dimension 3100 system (Bruker) to establish the thickness and homogeneity of the exfoliated 1T–TaS2 flakes. Line profiles were extracted across the flake edges from full-area scans. The resulting profiles (see Fig. S2), were leveled by subtracting the local slope and fitted to determine the step height, which corresponds to the flake thickness.
Scanning electron microscopy:
High-resolution images of the shadow mask were obtained using a scanning electron microscope (Regulus 8230, Hitachi).
Author contributions
@@
Data availability
Data of this article, including ARPES and AFM scans, transport data and photographs are available at the Zenodo repository under https://doi.org/10.5281/zenodo.15482370.
Conflict of Interest
@@ Acknowledgments
Programme, within Grant Agreement 810451 (HERO). R.V., A.M., Y.V and D.M. thank ARIS for funding the research program P1-0040, and D.M. thanks ARIS for funding the